# Almost Complete Intersections: Regularity of powers and Rees algebra

### v1

@@ 3.1. Almost complete intersections of grade $2$
Setup 2.
Let $B$ be a standard graded polynomial ring and $I=\langle f_{1},f_{2},f_{3}\rangle\subseteq B$ be an almost complete intersection of grade $2$ and perfect with deg $f_{i}=d_{i}$, $i=1,2,3$. Further assume that the presentation matrix of $I$ has the property that every non-zero entry in the $i^{th}$ row is of degree $\beta_{i}$.
@@
Consider the ideal $I$ as in Setup 2 and the notations as in Proposition 3.1. Then the minimal bigraded free resolution of the corresponding $\mathcal{R}(I)$ is of the following form:
$0\longrightarrow\begin{matrix}S(-(d_{1}+d_{2}+d_{3}),-2)\end{matrix}\stackrel{{\scriptstyle\delta_{2}}}{{\longrightarrow}}\begin{matrix}S(-(d_{k}+\beta_{k}),-1)\\ \oplus\\ S(-(d_{l}+\beta_{l}),-1)\end{matrix}\stackrel{{\scriptstyle\delta_{1}}}{{\longrightarrow}}S\longrightarrow 0$
where $\beta_{k}$ and $\beta_{l}$, $k,l\in\mathbb{N}$, corresponds to the degrees of elements from two different columns of the presentation matrix of $I$.
@@
From Hilbert-Burch theorem [12, Theorem 20.15], the graded minimal free resolution of $B/I$ has the form,
$0\longrightarrow\begin{matrix}B(-(d_{k}+\beta_{k}))\\ \oplus\\ B(-(d_{l}+\beta_{l}))\end{matrix}\stackrel{{\scriptstyle\phi_{2}}}{{\longrightarrow}}\begin{matrix}B(-d_{1})\\ \oplus\\ B(-d_{2})\\ \oplus\\ B(-d_{3})\end{matrix}\stackrel{{\scriptstyle\phi_{1}}}{{\longrightarrow}}B\longrightarrow 0$
where $\phi_{1}=\left[\begin{matrix}f_{1}&f_{2}&f_{3}\end{matrix}\right]$ and $\phi_{2}=\left[\begin{matrix}z_{11}&z_{12}\\ z_{21}&z_{22}\\ z_{31}&z_{32}\\ \end{matrix}\right]$ and further $\beta_{k}$ and $\beta_{l}$ come corresponding to some non-zero entries $z_{ks}$ and $z_{lt}$ respectively in $\phi_{2}$ such that $s-t\cong 1\mod(2)$.
@@
The result thus follows from Proposition 3.1. ∎
@@
Perfect ideals of grade $2$ are characterized by the Hilbert-Burch theorem [12]. Thus, any $3\times 2$ matrix $X$ with entries from a graded polynomial ring $B$ satisfying the condition that grade of the ideal of the corresponding maximal minors ($I_{2}(X)$) is bounded below by $2$ and, a non-zero divisor $z$ in $B$, corresponds to an ideal $zI_{2}(X)$ which has the form mentioned in Setup 2 and thus the associated Rees algebra is resolved using Proposition 3.6.
@@ 3.2. Almost complete intersections of grade $3$
Setup 3.
Let $B=K[X]$ denote a standard graded polynomial ring with $X$ being a skew-symmetric matrix of the form,
$X=\begin{bmatrix}0&0&0&x_{14}&x_{15}&\cdots&x_{1t}\\ 0&0&0&x_{24}&x_{25}&\cdots&x_{2t}\\ 0&0&0&x_{34}&x_{35}&\cdots&x_{3t}\\ -x_{14}&-x_{24}&-x_{34}&0&x_{45}&\cdots&x_{4t}\\ -x_{15}&-x_{25}&-x_{35}&-x_{45}&0&\cdots&x_{5t}\\ \vdots&\vdots&\vdots&\vdots&\vdots&\cdots&\vdots\\ -x_{1t}&-x_{2t}&-x_{3t}&-x_{4t}&-x_{5t}&\cdots&0\\ \end{bmatrix}$
of order $t$. Let $I=\langle f_{1},\ldots,f_{4}\rangle$ with deg $f_{i}=d_{i}$, $i=1,\ldots,4$, be an almost complete intersection ideal of $B$ of grade $3$ and type $t-3$ for $t\in\mathbb{N}$, $t\geq 5$.
@@
Consider the ideal $I$ as in Setup 3. Then the minimal bigraded free resolution of $\mathcal{R}(I)\cong S/J$ where $S=K[X,Y]$ with $Y=\begin{bmatrix}y_{1},\ldots,y_{4}\\ \end{bmatrix}$ and deg $x_{kl}=(1,0)$ and deg $y_{i}=(d_{i},1)$, $i=1,\ldots,4$ and $J$ is the defining ideal of $\mathcal{R}(I)$, has the following form:
@@
Let $I$ be an almost complete intersection of grade $3$ and type $t-3$. Following can be observed from the results in [7].
@@
Then the bigraded free resolutions of $\mathcal{R}(I)$ are consequences of Proposition 3.1.
@@
The bi-regularities of $\mathcal{R}(I)$ when $I$ is an almost complete intersection of grade $3$ of type $t-3$ are as follows,
@@ 3.3. Restriction on the degree of the generators of the ideal and certain colon ideals
    Assume $J^{\prime}=(J:f_{n+1})$ to be an equigenerated perfect ideal of grade $2$ with the degree of the generators being $d^{\prime}$, $\mu(J^{\prime})=t$, $t\in\mathbb{N}$ and $J^{\prime}$ being linearly presented. By Hilbert-Burch theorem, the graded minimal free resolution of $B/J^{\prime}$ has the following form:
@@
    As a consequence of Proposition 3.1, the following would be a (non-standard) bigraded free resolution of $\mathcal{R}(I)$,
@@
      The ideal $J^{\prime}$ need not always be equigenerated. For example let $I=\langle f_{1},f_{2},f_{3}\rangle$, an ideal in a polynomial ring $B$, be a complete intersection ideal. Then the defining relations of the corresponding Rees algebra is given by the maximal minors of the matrix $X=\begin{bmatrix}f_{1}&f_{2}&f_{3}\\ y_{1}&y_{2}&y_{3}\end{bmatrix}$. Let the relations be denoted by $\{g_{1},g_{2},g_{3}\}$. Then they satisfy the condition that $\{g_{1},g_{2}\}$ are $S$-regular and $(\langle g_{1},g_{2}\rangle:g_{3})=\langle f_{3},y_{3}\rangle$ ([6, Lemma 3.1]), which is clearly not equigenerated.
@@
      A class of ideals which satisfies the condition (a) are complete intersections minimally generated by $2$ linear terms. For example for $I=\langle x_{1}x_{2},x_{3}x_{4},x_{2}x_{3}\rangle\subset K[x_{1},x_{2},x_{3},x_{4}]$, then $\{x_{1}x_{2},x_{3}x_{4}\}$ forms a regular sequence and $(\langle x_{1}x_{2},x_{3}x_{4}\rangle:x_{2}x_{3})=\langle x_{1},x_{4}\rangle$ is a grade $2$ perfect ideal. In general, as mentioned in Remark 3.7, all perfect ideals of grade $2$ are characterized by the Hilbert Burch theorem.
@@
    From Proposition 3.1, the following would be a bigraded free resolution of $\mathcal{R}(I)$,
@@
In both the cases (a) and (b), $d^{\prime}\leq d$ since $J\subseteq J^{\prime}$. This needs to be kept in mind while studying the properties of diagonals in the next section.
@@ 4. Application to diagonal subalgebras of almost complete intersections
Consider the notations as in Proposition 3.1.
@@ 4.1. Koszulness of $\mathcal{R}(I)_{\Delta}$
Let $I=\langle f_{1},\ldots,f_{n+1}\rangle$ be an equigenerated ideal of $B$ satisfying the conditions in Proposition 3.1 with deg $f_{i}=d$. Then it is possible to associate a standard bigraded structure to the corresponding Rees algebra by considering it to be the quotient of a standard bigraded polynomial ring $S$ with deg $x_{i}=(1,0)$ and deg $y_{j}=(0,1)$ for $i=1,\ldots,m$ and $j=1,\ldots,n+1$. From the form of the differentials mentioned in Proposition 3.1, the standard bigraded free resolution of the Rees algebra will take the following form,
@@
It is clear from Complex (10) that the $y$-regularity of $\mathcal{R}(I)$ vanishes, and thus $\mathcal{R}(I)_{\Delta}$ is Koszul for $e>0$ [25, Theorem 2.9]. Thus only the $x$-shifts need to be looked into. Let $i$ denote the homological degree and $t_{i}^{B}(B/I)$ be as defined in Section 1. Then for $a_{max}^{i}=\text{max}\{t_{i+1}^{B}(B/I)+d,(i-1)d\}$, from Remark $(1)$, $\mathcal{R}(I)_{\Delta}$ is Koszul if $\dfrac{a_{max}^{i}}{c}\leq i+1$ for all $i\geq 1$.
Case 1: $a_{max}^{i}=(i-1)d$. Then for $c\geq d$, $\mathcal{R}(I)_{\Delta}$ is Koszul.
Case 2: $a_{max}^{i}=t_{i+1}^{B}(B/I)-d$. Then for $c\geq\dfrac{t_{i+1}^{B}(B/I)-d}{i+1}$, $\mathcal{R}(I)_{\Delta}$ is Koszul.
@@
As a consequence of the discussion above, following can be observed regarding the Koszulness of the diagonals of $\mathcal{R}(I)$ for the class of ideals discussed in Section 3.
@@
    Consider the assumptions as in Setup 2 with $d_{1}=d_{2}=d_{3}=d$, $d>1$ and $I$ being linearly presented. Then $\mathcal{R}(I)_{\Delta}$ is Koszul for $c\geq\frac{d}{3}$ and $e>0$.
@@
    Consider the assumptions as in Subsection 3.3$(a)$ and $(b)$. Then for $c\geq\big{(}\frac{n-1}{n}\big{)}d$ and $e>0$, $\mathcal{R}(I)_{\Delta}$ is Koszul.
@@
The previous proposition thus gives bounds on the diagonals of Rees algebra associated to almost complete intersections for which they are Koszul, similar to the ones given in [9, 6, 23] for Rees algebras corresponding to complete intersections.
@@ 4.2. Cohen-Macaulayness of $\mathcal{R}(I)_{\Delta}$
Let $I$ be a graded ideal of $B$ satisfying the conditions in any of the subsections 3.1, 3.2 or 3.3. Then $\mathcal{R}(I)$ is Cohen-Macaulay and $\mathcal{R}(I)_{\Delta}$ is Cohen-Macaulay for all $e>0$ and $c>\max\{\min\{(e-1)d+u-m,e(u-m)\},\min\{(e-1)d+u-d_{1},e(u-d_{1})\},ed\}$.
@@
Acknowledgement. The Core Research Grant (CRG/2023/007668) from the Science and Engineering Research Board, India, partially supports the first author. INSPIRE Fellowship, DST, India, financially supports the second author.

### v2

@@ 3.1. Almost complete intersections of grade $2$
Setup 3.
Let $B$ be a standard graded polynomial ring and $I=\langle f_{1},f_{2},f_{3}\rangle\subseteq B$ be an almost complete intersection of grade $2$ and perfect with deg $f_{i}=d_{i}$, $i=1,2,3$.
@@
Consider the ideal $I$ as in Setup 3 and the notations as in Theorem 3.1. Let $\beta_{ij}$ denote the degree of the $(i,j)^{th}$ element in the presentation matrix of $I$. Then the minimal bigraded free resolution of the corresponding $\mathcal{R}(I)$ is of the following form:
$0\longrightarrow\begin{matrix}S(-(d_{1}+d_{2}+d_{3}),-2)\end{matrix}\stackrel{{\scriptstyle\delta_{2}}}{{\longrightarrow}}\begin{matrix}S(-(d_{k}+\beta_{k1}),-1)\\ \oplus\\ S(-(d_{l}+\beta_{l2}),-1)\end{matrix}\stackrel{{\scriptstyle\delta_{1}}}{{\longrightarrow}}S\longrightarrow 0$
where $\beta_{k1}$ and $\beta_{l2}$ are non-zero.
@@
From the Hilbert-Burch theorem [13, Theorem 20.15], the graded minimal free resolution of $B/I$ has the form,
$0\longrightarrow\begin{matrix}B(-(d_{k}+\beta_{k1}))\\ \oplus\\ B(-(d_{l}+\beta_{l2}))\end{matrix}\stackrel{{\scriptstyle\phi_{2}}}{{\longrightarrow}}\begin{matrix}B(-d_{1})\\ \oplus\\ B(-d_{2})\\ \oplus\\ B(-d_{3})\end{matrix}\stackrel{{\scriptstyle\phi_{1}}}{{\longrightarrow}}B\longrightarrow 0$
where $\phi_{1}=\left[\begin{matrix}f_{1}&f_{2}&f_{3}\end{matrix}\right]$ and $\phi_{2}=\left[\begin{matrix}z_{11}&z_{12}\\ z_{21}&z_{22}\\ z_{31}&z_{32}\\ \end{matrix}\right]$ and $\beta_{k1}$ and $\beta_{l2}$ come corresponding to some non-zero entries $z_{k1}$ and $z_{l2}$ respectively in $\phi_{2}$.
@@
The result thus follows from Theorem 3.1. ∎
@@
Perfect ideals of grade $2$ are characterized by the Hilbert-Burch theorem [13]. Thus, any $3\times 2$ matrix $X$ with entries from a graded polynomial ring $B$ that satisfies the condition that the grade of the ideal of the corresponding maximal minors $I_{2}(X)$ is bounded below by $2$, along with a non-zero divisor $z$ in $B$, corresponds to an ideal $zI_{2}(X)$, which has the form mentioned in Setup 3. Consequently, the associated Rees algebra is resolved using Proposition 3.6.
@@ 3.2. Almost complete intersections of grade $3$
Consider the ideal $I$ as in Setup 2. Then the minimal bigraded free resolution of $\mathcal{R}(I)\cong S/J$ where $S=K[X,Y]$ with $Y=\begin{bmatrix}y_{1},\ldots,y_{4}\\ \end{bmatrix}$ and deg $x_{kl}=(1,0)$ and deg $y_{i}=(d_{i},1)$, $i=1,\ldots,4$ and $J$ is the defining ideal of $\mathcal{R}(I)$, has the following form:
@@
Let $I$ be an almost complete intersection of grade $3$ and type $t-3$. The following can be observed from the results in [8].
@@
Then the bigraded free resolutions of $\mathcal{R}(I)$ are consequences of Theorem 3.1.
@@
The bi-regularities of $\mathcal{R}(I)$ when $I$ is an almost complete intersection of grade $3$ of type $t-3$ in a standard graded polynomial ring are as follows,
@@ 3.3. Restriction on the degree of the generators of the ideal and certain colon ideals
    Assume $J^{\prime}=(J:f_{n+1})$ to be an equigenerated perfect ideal of grade $2$ with the degree of the generators being $d^{\prime}$, $\mu(J^{\prime})=t$, $t\in\mathbb{N}$ and $J^{\prime}$ being linearly presented. By the Hilbert-Burch theorem, the graded minimal free resolution of $B/J^{\prime}$ has the following form:
@@
    As a consequence of Theorem 3.1, the following would be a (non-standard) bigraded free resolution of $\mathcal{R}(I)$,
@@
      The ideal $J^{\prime}$ need not always be equigenerated. For example let $I=\langle f_{1},f_{2},f_{3}\rangle$, an ideal in a polynomial ring $B$, be a complete intersection ideal. Then the defining relations of the corresponding Rees algebra is given by the maximal minors of the matrix $X=\begin{bmatrix}f_{1}&f_{2}&f_{3}\\ y_{1}&y_{2}&y_{3}\end{bmatrix}$. Let the relations be denoted by $\{g_{1},g_{2},g_{3}\}$. Then they satisfy the condition that $\{g_{1},g_{2}\}$ are $S$-regular and $(\langle g_{1},g_{2}\rangle:g_{3})=\langle f_{3},y_{3}\rangle$ ([7, Lemma 3.1]), which is clearly not equigenerated.
@@
      An example of a class of ideals $J^{\prime}$ that satisfies the condition (A) are complete intersections minimally generated by $2$ linear forms. For example, for $I=\langle x_{1}x_{2},x_{3}x_{4},x_{2}x_{3}\rangle\subset K[x_{1},x_{2},x_{3},x_{4}]$, the set $\{x_{1}x_{2},x_{3}x_{4}\}$ forms a regular sequence, and $(\langle x_{1}x_{2},x_{3}x_{4}\rangle:x_{2}x_{3})=\langle x_{1},x_{4}\rangle$ is a grade $2$ perfect ideal. In general, as mentioned in Remark 3.7, all perfect ideals of grade $2$ are characterized by the Hilbert Burch theorem.
@@
    From Theorem 3.1, the following would be a bigraded free resolution of $\mathcal{R}(I)$,
@@
In both cases (A) and (B), we have $d^{\prime}\leq d$ since $J\subseteq J^{\prime}$. This should be kept in mind while studying the properties of diagonals in the next section.
@@ 4. Application to diagonal subalgebras of almost complete intersections
Consider the notations as in Theorem 3.1.
@@ 4.1. Koszulness of $\mathcal{R}(I)_{\Delta}$
Let $I=\langle f_{1},\ldots,f_{n+1}\rangle$ be an equigenerated ideal of $B$ satisfying the conditions in Theorem 3.1 with deg $f_{i}=d$. Then it is possible to associate a standard bigraded structure to the corresponding Rees algebra by considering it to be the quotient of a standard bigraded polynomial ring $S$ with deg $x_{i}=(1,0)$ and deg $y_{j}=(0,1)$ for $i=1,\ldots,m$ and $j=1,\ldots,n+1$. From the form of the differentials mentioned in Theorem 3.1, the standard bigraded free resolution of the Rees algebra will take the following form,
@@
Let $i$ denote the homological degree and $t_{i}^{B}(B/I)$ be as defined in Section 1. Let $a_{\max}^{i}$ and $b_{\max}^{i}$ denote the maximum shifts in the first and second degrees, respectively, at the $i^{th}$ homological degree of the Complex (19). Clearly,
$a_{\max}^{i}=\max\{t_{i+1}^{B}(B/I)+d,(i-1)d\}$
and
$b_{\max}^{i}=\max\{j+2\mid j=0,\ldots,i-1\}=i+1.$
Then from Remark 1.4(1), $\mathcal{R}(I)_{\Delta}$ is Koszul if $\dfrac{a_{max}^{i}}{c}\leq i+1$ and $\dfrac{b_{\max}^{i}}{e}\leq i+1$ for all $i\geq 1$.
Since $b_{\max}^{i}=i+1$, the condition $\dfrac{b_{\max}^{i}}{e}\leq i+1$ holds true for all $e>0$. It remains to check the values of $c$ for which $\dfrac{a_{\max}^{i}}{c}\leq i+1.$ For this, we need to consider the following two cases from the Complex (19),
Case 1: $a_{max}^{i}=(i-1)d$. Then $\dfrac{a_{max}^{i}}{c}\leq i+1\Longleftrightarrow\dfrac{(i-1)d}{c}\leq i+1\Longleftrightarrow\dfrac{i-1}{i+1}d\leq c\Longleftrightarrow c\geq d$ since $\left\lvert{\dfrac{i-1}{i+1}}\right\rvert\leq 1$. Thus for $c\geq d$ and $e>0$, $\mathcal{R}(I)_{\Delta}$ is Koszul.
Case 2: $a_{max}^{i}=t_{i+1}^{B}(B/I)-d$. Then for $c\geq\dfrac{t_{i+1}^{B}(B/I)-d}{i+1}$ and $e>0$, $\mathcal{R}(I)_{\Delta}$ is Koszul.
@@
As a consequence of the discussion above, the following can be observed regarding the Koszulness of the diagonals of $\mathcal{R}(I)$ for the class of ideals discussed in Section 3.
@@
    Consider the assumptions as in Setup 3 with $d_{1}=d_{2}=d_{3}=d$, $d>1$ and $I$ being linearly presented. Then $\mathcal{R}(I)_{\Delta}$ is Koszul for $c\geq\frac{d}{3}$ and $e>0$.
@@
    Let $I$ be an equigenerated ideal of height $n$ with a generating set $\{f_{1},\ldots,f_{n+1}\}$ such that $f_{1},\ldots,f_{n}$ forms a regular sequence. Further, let $J=\langle f_{1},\ldots,f_{n}\rangle$ and $J^{\prime}=(J:f_{n+1})$ satisfy the assumptions as in Subsection 3.3 (A) or (B). Then for $c\geq\big{(}\frac{n-1}{n}\big{)}d$ and $e>0$, $\mathcal{R}(I)_{\Delta}$ is Koszul.
@@
Note that Proposition 4.1 provides bounds on the diagonals of the Rees algebra associated with almost complete intersections, for which they are Koszul, similar to those found for the Rees algebra corresponding to complete intersections [10, 7, 24], residual intersections [1] and maximal order Pfaffians [25].
@@ 4.2. Cohen-Macaulayness of $\mathcal{R}(I)_{\Delta}$
Let $I=\langle f_{1},\ldots,f_{n+1}\rangle$ be a graded ideal in a standard graded polynomial ring $B=K[x_{1},\ldots,x_{m}]$ with the degree of $f_{i}=d_{i}$, satisfying either of the following conditions:
  1. (1)
    $I$ be an almost complete intersection ideal satisfying the assumptions as in Setup 2 or Setup 3.
  2. (2)
    The height of the ideal $I$ is $n$ with $f_{1},\ldots,f_{n}$ forming a regular sequence. Further, let $J=\langle f_{1},\ldots,f_{n}\rangle$ and $J^{\prime}=(J:f_{n+1})$ satisfy the assumptions as in Subsection 3.3 (A) or (B).
Set $d=\max\{d_{1},\ldots,d_{n+1}\}$ and $u=\sum_{i=1}^{n+1}d_{i}$. Then $\mathcal{R}(I)$ is Cohen-Macaulay and $\mathcal{R}(I)_{\Delta}$ is Cohen-Macaulay for all $e>0$ and $c>\max\{\min\{(e-1)d+u-m,e(u-m)\},\min\{(e-1)d+u-d_{1},e(u-d_{1})\},ed\}$.
@@
Acknowledgement. The Core Research Grant (CRG/2023/007668) from the Science and Engineering Research Board, ANRF, India, partially supports the first author. INSPIRE Fellowship, DST, India, financially supports the second author.